OpenStudy (anonymous):

Karl and Patrick are playing with a fair coin. The one who obtains more tails than heads in 4 tosses is declared the winner. What is the probability that Patrick wins?

OpenStudy (nurali):

these are tricky counting problems where one has to be careful ... if they each toss the coin 4 times, each guy will have 16 possible outcomes (2^4) and the prob. of a any specific outcome is 1/16. there is 1 way of getting all (4) tails, 4 ways of getting 3 tails, 6 ways of 2 tails, 4 ways of 1 tail, and 1 way of no tails (ie all heads). so the probs are: P(0 tails)=1/16 P(1 tail)=4/16 P(2 tails)=6/16 (not 1/2 !) P(3 tails)=4/16 P(4 tails)=1/16 we need P(# of Patrick tails > # of Karl tails) = P(1T and 0T)+P(2T and (0 or 1) T)+P(3T and (0 or 1 or 2) T)+P(4T and (0 or 1 or 2 or 3) T) = (4/16*1/16)+(6/16*5/16)+(4/16*11/16)+(1/16*15/16) = 93/256 = 0.36328 another way of lookin at it: The prob that Patrick wins = prob that Karl wins = (1 - P(they tie)) / 2 prob they tie = P(# of Patrick tails = # of Karl tails) = (1/16*1/16)+(4/16*4/16)+(6/16*6/16)+(4/16*4/16)+(1/16*1/16) = 1/256 + 16/256 + 36/256 + 16/256 + 1/256 = 70/256 = 0.2734 therefore, to check... P(Patrick wins) or P(Karl wins) or P(they tie) = 93/256 + 93/256 + 70/256 = 1 as the probability of all possible outcomes must total 1